A family has two cars. During one particular week, the first car consumed 20 gallons of gas and the second consumed 25 gallons of gas. The two cars drove a combined total of 1375 miles, and the sum of their fuel efficiencies was 60 miles per gallon. What were the fuel efficiencies of each of the cars that week
First car = how many gallons?
Second car = how many gallons?

Found 2 solutions by Cromlix, prekteacherlisa:
Answer by Cromlix(4381)

First car = x
Second car = y
x + y = 55.......1
35x + 25y = 1675.....2
Multiply (1) by 25
25x + 25y = 1375...3
35x + 25y = 1675...2
Subtract (3) from (2)
10x = 300
x = 30
Substitute x = 30 into (1)
x + y = 55
30 + y = 55
y = 55 - 30
y = 25
First car 30 gallons
Second car 25 gallons

Answer by prekteacherlisa(1)

x+y = 60
20x + 25y=1375
25x +25y=1500
5x=125
x=25 car 1
25 + y=60
y=60-25
y=35 car 2